OpenStudy (anonymous):

what's the geometric mean of 31 and 42? do I just multiply them and then square root?

jimthompson5910 (jim_thompson5910):

you are absolutely correct

OpenStudy (anonymous):

after i multiply them my teacher told us to find two #'s that you can multiply and get that product, but one of those has to be a perfect square :( so i'm stuck at 1302 because what number can i multiply to get that but one has to be a perfect square ....

jimthompson5910 (jim_thompson5910):

try listing the list of perfect squares 1, 4, 9, 16, 25, 36, ... and see which number is a factor of 1302

jimthompson5910 (jim_thompson5910):

it turns out that 1302 has no such factors we can prove this by looking at the prime factorization of 1302

jimthompson5910 (jim_thompson5910):

1302 = 31*42 1302 = 31*6*7 1302 = 31*2*3*7 Notice we have no repeat factors. So that means there are no perfect square factors

OpenStudy (anonymous):

I will leave it as that then, thank you! :)

jimthompson5910 (jim_thompson5910):

you're welcome